# Federated LLMs Fine-tuned with Adaptive Importance-Aware LoRA

### v1

@@ III-D Adaptive Global Aggregation at the Cloud
In the HAFL framework, after completing local training, each client uploads its updated LoRA parameters to the cloud server. These parameters are then aggregated to form a new global model. For the truncation-based scheme, the upload LoRA parameters are $\mathbf{B}_{k}$ and $\mathbf{A}_{k}$. In contrast, for the importance-based parameter freezing scheme, the upload LoRA parameters are $\mathbf{B}_{k}^{\mathcal{I}_{k}}=\mathbf{B}_{k}[:,\mathcal{I}_{k}]$ and $\mathbf{A}_{k}^{\mathcal{I}_{k}}=\mathbf{A}_{k}[\mathcal{I}_{k},:]$.
Adaptive Aggregation of LoRA Parameters
The adaptive aggregation is shown in Algorithm 1, rather than aggregating LoRA parameters at the scale of the entire LoRA matrices, we perform aggregation at the decomposed rank-1 matrix level. For each rank-1 matrix, only the clients that update and upload that rank-1 matrix will participate in the aggregation process. Each client $k$ contributes to the global model based on a sparsity-based norm $z_{k}$, defined as
$z_{k}=\begin{cases}\|\mathbf{B}_{k}\mathbf{A}_{k}\|_{F},&\text{Truncation scheme}\\ \|\mathbf{B}_{k}^{\mathcal{I}_{k}}\mathbf{A}_{k}^{\mathcal{I}_{k}}\|_{F},&\text{Freezing scheme}\end{cases},$ (14)
@@
0:  LoRA parameters $\mathbf{B}_{k},\mathbf{A}_{k}$ for each client $k\in\mathcal{K}$ in parameter truncation scheme or LoRA parameters $\mathbf{B}_{k}^{\mathcal{I}_{k}},\mathbf{A}_{k}^{\mathcal{I}_{k}}$ in parameter freezing scheme.
@@
11:           $\mathbf{B}_{g}^{(t+1)}[:,j]\mathrel{+}=z_{k}\mathrel{/}Z[j]*\mathbf{B}_{k}[:,i]$
12:           $\mathbf{A}_{g}^{(t+1)}[j,:]\mathrel{+}=z_{k}\mathrel{/}Z[j]*\mathbf{A}_{k}[i,:]$
@@
14:           $\mathbf{B}_{g}^{(t+1)}[:,j]\mathrel{+}=z_{k}\mathrel{/}Z[j]*\mathbf{B}_{k}^{\mathcal{I}_{k}}[:,i]$
15:           $\mathbf{A}_{g}^{(t+1)}[j,:]\mathrel{+}=z_{k}\mathrel{/}Z[j]*\mathbf{A}_{k}^{\mathcal{I}_{k}}[i,:]$
@@
IV Numerical Results
In our experiments, we utilize GPT2-Large[12] as the backbone model and test on the 20 News Group classification dataset[13]. The experiment involves 100 clients and one server, with the dataset evenly distributed among these 100 clients. We assume that different clients have diverse computational capabilities, which correspond to different LoRA ranks. Specifically, in the parameter truncation scheme, 33 clients have a LoRA rank of 2, another 33 clients have a LoRA rank of 4, and the remaining 34 clients have a LoRA rank of 16. The global model at the server is configured with a LoRA rank of 16. In the parameter freezing scheme, all clients and the cloud server maintain a uniform LoRA rank of 16. However, 33 clients have a freezing ratio of 0.875, another 33 clients have a freezing ratio of 0.75, and the remaining 34 clients have a freezing ratio of 0. To simulate a real-world FL setting, 10 clients are randomly selected to participate in the training during each communication round. The hyper-parameters related to the experiment are shown in Table I.
TABLE I: Experiment Hyper-parameters
@@ IV Numerical Results
Figure 2: Accuracy comparison on 20 News Group Classification
Fig. 2 presents the classification accuracy of the global model on the 20 News Group classification test set as the number of communication rounds increases. Our methods, Importance-based Parameters Freezing Adaptive Aggregation LoRA (IFALoRA) and Importance-based Parameters Truncation Adaptive Aggregation LoRA (ITALoRA), are shown with fast convergence. We also compared three other algorithms: (1) Homogeneous LoRA (HomLoRA) $r=2$, where all clients and cloud server use LoRA with a rank of 2; (2) HomLoRA $r=16$, where all clients and cloud server use LoRA with a rank of 16; and (3) IFZLoRA, which utilizes zero-padding aggregation compared to IFALoRA. The results indicate that HomLoRA $r=2$ fails to converge, and IFZLoRA converges slowly with significantly lower accuracy than our TrALoRA and IFALoRA. Our methods achieve accuracy close to that of HomLoRA $r=16$, where all clients use the highest rank.
TABLE II: Evaluation results on 20 News Group classification
Model Method Commu. Size 20NEWS(Acc.)
@@
GPT2-Large IFZLoRA 52.28M 36.43±7.6 62.83±1.3
HomLoRA r=16 113M 67.41±0.6 69.74±0.2
HomLoRA r=2 14M 6.25±1.0 5.66±0.7
ITALoRA 52.28M 65.30±1.8 67.31±0.3
IFALoRA 52.28M 65.93±0.8 67.55±0.8
@@
The Table II presents the performance of five different methods on the 20 News text classification task after 50 and 100 communication rounds. The values are averaged over experiments with different random seeds, and the table also records the average total parameter size uploaded by 10 clients per communication round. Specifically, HomLoRA with rank 16 achieves the highest accuracy but requires the largest communication size. In contrast, IFALoRA and ITALoRA manage to achieve nearly comparable accuracy with significantly smaller communication sizes, demonstrating a good balance between performance and efficiency. HomLoRA with rank 2 shows poor performance, indicating a lack of convergence, while IFZLoRA improves over HomLoRA r = 2 but still lags behind ITALoRA and IFALoRA in terms of accuracy.
@@
Fig. LABEL:fig:Performance_degradation LABEL:fig:ITALoRA shows the performance degradation of the ITALoRA method when the global model is truncated to smaller client ranks. As the truncation increases, the model’s accuracy decreases, indicating a loss of information. In contrast, our IFALoRA method in Fig. LABEL:fig:Performance_degradation LABEL:fig:IFALoRA uses a freezing strategy, eliminating the need for truncation. This allows each client to maintain the same accuracy performance level as the global model. Additionally, both IFALoRA and ITALoRA have the same parameter size during training, ensuring communication efficiency while IFALoRA avoids the performance loss associated with truncation.
V Conclusion
@@
In this paper, we proposed a HAFL framework to address privacy concerns in fine-tuning LLMs. To accommodate client resource heterogeneity, we first introduced an importance-based parameter truncation scheme at the local LLM. However, the truncation process inevitably results in performance degradation. To tackle this, we further developed an importance-based parameter frozen scheme at the local LLM. At the cloud server, we proposed an adaptive global model aggregation method to counter the information dilution problem caused by the zero-padding aggregation method. Our experimental results demonstrate that our method achieves rapid convergence with low communication overhead. Compared to the truncation approach, our importance-based partial freezing scheme maintains client model performance during global model parameter distribution. Furthermore, the adaptive aggregation method achieves faster convergence compared to the zero-padding approach.

### v2

@@ VI Adaptive Global Aggregation at the Cloud
After dequantization, the server collects the uploaded LoRA parameters from all clients and aggregates them to form a new global model. For the truncation-based scheme, the upload dequantized LoRA parameters are $\mathbf{B}_{k,\text{dq}}^{\text{T}}$ and $\mathbf{A}_{k,\text{dq}}^{\text{T}}$. In contrast, for the importance-based parameter freezing scheme, the upload dequantized LoRA parameters are $\mathbf{B}_{k,\text{dq}}^{\mathcal{I}}$ and $\mathbf{A}_{k,\text{dq}}^{\mathcal{I}}$.
@@
0:  Dequantized LoRA parameters $\mathbf{B}_{k,\text{dq}}^{\text{T}},\mathbf{A}_{k,\text{dq}}^{\text{T}}$ for each client $k\in\mathcal{K}$ in parameter truncation scheme or Dequantized LoRA parameters $\mathbf{B}_{k,\text{dq}}^{\mathcal{I}},\mathbf{A}_{k,\text{dq}}^{\mathcal{I}}$ in parameter freezing scheme.
@@
11:           $\mathbf{B}_{g}^{(t+1)}[:,j]\mathrel{+}=z_{k}\mathrel{/}Z[j]\cdot\mathbf{B}_{k,\text{dq}}^{\text{T}}[:,i]$
12:           $\mathbf{A}_{g}^{(t+1)}[j,:]\mathrel{+}=z_{k}\mathrel{/}Z[j]\cdot\mathbf{A}_{k,\text{dq}}^{\text{T}}[i,:]$
@@
14:           $\mathbf{B}_{g}^{(t+1)}[:,j]\mathrel{+}=z_{k}\mathrel{/}Z[j]\cdot\mathbf{B}_{k,\text{dq}}^{\mathcal{I}}[:,i]$
15:           $\mathbf{A}_{g}^{(t+1)}[j,:]\mathrel{+}=z_{k}\mathrel{/}Z[j]\cdot\mathbf{A}_{k,\text{dq}}^{\mathcal{I}}[i,:]$
@@
Adaptive Aggregation of LoRA Parameters
The adaptive aggregation is shown in Algorithm 2. Rather than aggregating LoRA parameters at the scale of the entire LoRA matrices, we perform aggregation at the decomposed rank-1 matrix level. For each rank-1 matrix, only the clients that update and upload that rank-1 matrix will participate in the aggregation process. Each client $k$ contributes to the global model based on a sparsity-based norm $z_{k}$, defined as
$z_{k}=\begin{cases}\|\mathbf{B}_{k,\text{dq}}^{\text{T}}\mathbf{A}_{k,\text{dq}}^{\text{T}}\|_{F},&\text{Truncation scheme}\\ \|\mathbf{B}_{k,\text{dq}}^{\mathcal{I}}\mathbf{A}_{k,\text{dq}}^{\mathcal{I}}\|_{F},&\text{Freezing scheme}\end{cases},$ (38)
VII Numerical Results
In our experiments, we utilize GPT2[30] as the backbone model and test on the Banking77 classification dataset[31]. The experiment involves 10 clients, one base station, and one server, with the dataset evenly distributed among these 10 clients. We assume that different clients have diverse computational capabilities, which correspond to different LoRA ranks or freezing ratios. Specifically, in the parameter truncation scheme, clients 0 to 2 have a LoRA rank of 2, clients 3 to 5 have a LoRA rank of 4, and clients 6 to 9 have a LoRA rank of 8. The global model at the server is configured with a LoRA rank of 8. In the parameter freezing scheme, all clients and the cloud server maintain a uniform LoRA rank of 8. However, clients 0 to 2 have a freezing ratio of 0.75, clients 3 to 5 have a freezing ratio of 0.5, and clients 6 to 9 have a freezing ratio of 0. All 10 clients participate in the training during each communication round. Additionally, the clients are positioned at increasing distances from the base station, ranging from 1100 meters to 2000 meters in increments of 100 meters. Specifically, client 0 is located 1100 meters away, client 1 at 1200 meters, and so forth, with client 9 positioned at 2000 meters. The hyper-parameters for the experiments are shown in Table I, while the wireless simulation parameters are detailed in Table II.
TABLE I: Experiment hyper-parameters.
@@ VII Numerical Results
TABLE II: Wireless system parameters.
Parameter Value Parameter Value
Carrier Freq. $f$ 2.4 GHz Bandwidth $B_{0}$ 100, 10 MHz
Commu. Time $T_{\text{ul}}$ 10 ms UE Tx Power $P_{k}$ 23 dBm
PSD $N_{0}^{\text{bs}}$174 dBm/Hz Std. Dev. $\sigma$ 7.8 dB
Figure 2: GPU memory usage for clients with different computing capabilities.
Fig. 2 plots the GPU memory usage for clients with varying computational capabilities under our salience-driven adaptive LLM quantization algorithm. The GPT-2 model used in this experiment consists of 12 transformer blocks. For clients with low computational capabilities, all transformer blocks are quantized to 8 bits, resulting in the lowest GPU memory usage. For clients with medium computational capabilities, the algorithm ranks transformer blocks by salience, keeping 32-bit precision for the 6 most salient blocks while quantizing the remaining 6 blocks to 8 bits, achieving a balance between memory efficiency and model precision. For clients with high computational capabilities, all transformer blocks keep 32-bit precision, leading to the highest GPU memory usage. These results demonstrate the effectiveness of our proposed algorithm in dynamically adjusting the quantization strategy to optimize resource utilization while accommodating the varying computational capabilities of different clients.
Figure 3: Accuracy comparison on Banking77 classification.
Fig. 3 presents the classification accuracy of the global model on the Banking77 classification test set as the number of communication rounds increases. In this experiment, we set the available bandwidth for each client to $B_{0}=100$ MHz, ensuring that communication bandwidth is not a limiting factor for FL parameter uploads. Our methods, importance-based parameter freezing adaptive aggregation LoRA (IFALoRA) and importance-based parameter truncation adaptive aggregation LoRA (ITALoRA), yield fast convergence. We also compared with three other algorithms: (1) homogeneous LoRA (HomLoRA) $r=2$, where all clients and the cloud server use LoRA with a rank of 2; (2) HomLoRA $r=8$, where all clients and the cloud server use LoRA with a rank of 8; and (3) IFZLoRA, which utilizes zero-padding aggregation compared to IFALoRA. The results indicate that HomLoRA $r=2$ fails to converge, and IFZLoRA converges slowly with significantly lower accuracy than our ITALoRA and IFALoRA. Our methods achieve accuracies close to that of HomLoRA $r=8$, where all clients use the highest rank.
TABLE III: Evaluation results on Banking77 classification.
Model Method Commu. Size Classification Accuracy
@@
GPT2 IFZLoRA 7.03M 71.34±2.4 82.06±1.2
HomLoRA r=8 11.25M 87.48±0.0 89.94±0.2
HomLoRA r=2 2.81M 21.26±14.0 45.79±16.0
ITALoRA 7.03M 82.53±2.1 86.70±0.5
IFALoRA 7.03M 87.36±0.2 89.13±0.2
@@
Table III presents the performance of five different methods on the Banking77 text classification task after 50 and 100 communication rounds. The values are averaged over experiments with different random seeds, and the table also records the average total parameter size uploaded by 10 clients per communication round. Specifically, HomLoRA with rank 8 achieves the highest accuracy but requires the largest communication size. In contrast, IFALoRA and ITALoRA manage to achieve nearly comparable accuracy with significantly smaller communication sizes, demonstrating a good balance between performance and efficiency. HomLoRA with rank 2 shows poor performance, indicating a lack of convergence, while IFZLoRA improves over HomLoRA r = 2 but still lags behind ITALoRA and IFALoRA in terms of accuracy.
@@
Fig. LABEL:fig:Performance_degradation LABEL:fig:ITALoRA plots the classification accuracy with the increasing of communication rounds under ITALoRA when the global model is truncated to smaller client ranks. As the level of truncation increases, the model’s accuracy decreases, indicating a loss of information. In contrast, our IFALoRA method in Fig. LABEL:fig:Performance_degradation LABEL:fig:IFALoRA uses a freezing strategy, eliminating the need for truncation. This allows each client to maintain the same accuracy performance level as the global model. Additionally, both IFALoRA and ITALoRA have the same parameter size during training, ensuring communication efficiency while IFALoRA avoids the performance loss associated with truncation.
Figure 5: Maximum number of transmission bits and computational capability.
Fig. 5 depicts the distribution of the maximum number of transmission bits and computational capability among different clients. The red bars represent the maximum number of transmission bits, while the purple bars indicate the computational capability levels. With a fixed bandwidth allocation of $B_{0}=10$ MHz for each client, the maximum number of transmission bits vary based on the distance to the base station. Client 0, being the closest to the base station, can support the highest number of transmission bits. As the distance increases, the maximum number of transmission bits gradually decreases, with client 9 having the lowest value. For computational capability, we assign different levels to clients: clients 0 to 2 have the lowest computational capability, clients 3 to 5 are assigned a medium computational capability, and clients 6 to 9 have the highest computational capability.
Figure 6: Importance-aware bandwidth-adaptive communication quantization scheme.
Fig. 6 plots the post-quantization communication size and the size reduction achieved for each client against the maximum transmission bits under the proposed importance-aware bandwidth-adaptive communication quantization scheme. Clients 0 to 2, which are relatively closer to the base station, benefit from higher maximum transmission bits and can upload their full-precision LoRA parameters without requiring quantization. In contrast, clients 3 to 9 have to be quantized to reduce the original LORA parameters’ size under the limit of maximum transmission bits via our proposed adaptive quantization algorithm.
Figure 7: Communication Quantization distribution of importance-aware bandwidth-adaptive communication quantization.
Fig. 7 presents the communication quantization distribution across different clients under our importance-aware bandwidth-adaptive communication quantization algorithm. The quantization levels are dynamically adjusted based on the clients’ communication conditions, leading to the following trends: Clients 0 to 2 are positioned close to the base station, benefiting from higher maximum transmission bits. As a result, they exclusively use 32-bit precision for transmission. Clients 3 to 5 adopt a mixed-precision scheme combining 32-bit and 16-bit quantization. As the distance from the base station increases, the proportion of 16-bit precision gradually rises, reflecting the need for reduced communication overhead. Clients 6 and 7 employ a mix of 16-bit and 8-bit quantization, further lowering communication costs while maintaining performance. Clients 8 and 9 utilize the most aggressive quantization strategy, relying on a combination of 8-bit and 4-bit precision to stay within bandwidth constraints.
Figure 8: Accuracy comparison of different communication precision schemes.
Fig. 8 plots the classification accuracy comparison of different communication quantization schemes. All considered schemes are based on the IFALoRA framework, with different communication precision methods: full-precision (32-bit), half-precision (16-bit), 8-bit quantization, and 4-bit quantization. These methods maintain a fixed precision throughout training, prioritizing the transmission of important LoRA parameters first. If the number of parameters exceeds the bandwidth limit, the remaining ones are discarded. Among all schemes, our proposed IFALoRA with importance-aware bandwidth-adaptive quantization scheme achieves the fastest convergence and the highest final accuracy. By dynamically adjusting the precision of the LoRA parameters based on bandwidth constraints, it ensures efficient communication while preserving model performance. Among the fixed-precision quantization methods, the 4-bit quantization scheme achieves the fastest initial convergence due to its low communication overhead, which allows more parameters to be transmitted within the bandwidth limit. However, the high quantization error in 4-bit precision accumulates over communication rounds, leading to noisier updates and the lowest final accuracy. The 32-bit full-precision scheme, on the other hand, has the slowest convergence because of its high communication cost, which limits the number of parameter updates within the same budget. Nevertheless, it achieves relatively higher final accuracy by preserving all parameter information. The 16-bit and 8-bit quantization methods offer a good trade-off, achieving faster convergence than full precision while resulting in minimal loss in final accuracy.
Figure 9: Average uplink data size of different communication precision schemes.
Fig. 9 presents a comparison of our importance-aware bandwidth adaptive communication quantization scheme against fixed-precision quantization methods in terms of the average uplink data size per client. The figure illustrates four key components: the original data size, the quantized data size, the actual transmitted data, and the discarded data due to bandwidth constraints. Our adaptive quantization scheme ensures that the quantized data size precisely aligns with the available bandwidth, thereby eliminating any discarded data caused by bandwidth limitations. In contrast, the 32-bit full-precision scheme transmits significantly more data than the bandwidth can accommodate, resulting in nearly half of the data being discarded. As the bit width decreases, the proportion of discarded data gradually reduces, with the 16-bit and 8-bit quantization schemes exhibiting progressively lower levels of discarded data. The 4-bit quantization scheme remains entirely within the bandwidth limit, leading to no discarded data. From a channel utilization perspective, both our adaptive scheme and the 32-bit full-precision scheme fully utilize the available bandwidth. However, while the 32-bit scheme results in substantial discarded data due to bandwidth constraints, our adaptive approach achieves optimal channel utilization by ensuring that all transmitted data fits within the available bandwidth, avoiding unnecessary data discard. In contrast, lower-precision schemes such as 4-bit quantization, despite preventing discarded data, do not fully exploit the available bandwidth, leading to reduced channel utilization.
Figure 10: Comparison of performance metrics with incorporation of adaptive schemes.
Fig. 10 plots a comprehensive performance comparison when progressively incorporating adaptive aggregation, adaptive communication quantization, and adaptive LLM quantization, all built upon the foundation of the importance-based parameter freezing scheme, evaluated in terms of model accuracy, discarded communication bits, and GPU memory usage. First, adaptive aggregation significantly improves accuracy from 59.38% (under ZeroPadding aggregation) to 83.60% by mitigating parameter dilution. Next, incorporating adaptive communication quantization further boosts accuracy to 88.73% by dynamically adjusting quantization levels in response to bandwidth constraints, thereby eliminating discarded bits (reduced from 2.88 Mbits to 0). Finally, the inclusion of adaptive LLM quantization reduces GPU memory usage from 519.13 MB to 360.12 MB while maintaining high accuracy (88.89%), thereby demonstrating its effectiveness in optimizing resource consumption without compromising performance.
VIII Conclusions
@@
In this paper, we proposed the Heterogeneous Adaptive Federated LoRA Fine-tuned LLM with Quantization framework to address privacy concerns in fine-tuning LLMs. To accommodate client resource heterogeneity, we first designed a salience-driven adaptive quantization scheme that dynamically adjusts the proportion of quantized transformer blocks based on both the computational capabilities of the clients and the importance of the blocks. To further address resource heterogeneity, we introduced an importance-based parameter truncation scheme at the local LLM. However, the truncation process inevitably resulted in performance degradation. To mitigate this, we developed an importance-based parameter freezing scheme at the local LLM. For the cloud server, we proposed an adaptive global model aggregation method to counter the information dilution problem caused by the zero-padding aggregation method. Moreover, we proposed an Importance-Aware Bandwidth-Adaptive Communication Quantization scheme, which prioritized the transmission of important LoRA rank-1 matrices with higher precision under bandwidth constraints. Experimental results demonstrated the effectiveness of our framework, with GPU memory usage decreasing by 31%, communication cost reduced by 49%, and model accuracy improving by 50%. These results highlights that our framework effectively supports federated fine-tuning of LLMs in real-world heterogeneous environments.